OpenStudy (anonymous):

if the slope of line l is -2/5, what is the slope of line m so that line l is perpendicular to line m? a. -2/5 b. -5/2 c. 2/5 d. 5/2

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (kagıtucak):

their slope multiplation must be -1

OpenStudy (anonymous):

i think it is c

OpenStudy (kagıtucak):

m.(-2/5)=-1

OpenStudy (kagıtucak):

m=5/2

OpenStudy (anonymous):

so its D?

OpenStudy (kagıtucak):

yea D

OpenStudy (anonymous):

here is the equation to get the slope of a perpendicular line: \[m_1(m_2)=-1\] let m_1 represent the slope of the line we know, -2/5 and m_2 represents the slope of the line that we want to be perpendicular. plug in m_1 = -2/5 into the equation: \[(\frac {-2}{5})(m_2)=-1\] solve for m_2